OpenStudy (anonymous):

i need help i will give you a metal for the answer plz Find a rational number and an irrational number that are between 7.7 and 7.9. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

Rational numbers can be expressed as fractions.

OpenStudy (anonymous):

7 8/10 is one. It simplifies to 7.8

OpenStudy (anonymous):

ok thanks can you help me on one more

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

Annabelle and Navene are multiplying (4275)(4372). Annabelle's Work Navene's Work (4275)(4372) = 42 + 375 + 2 = 4577 (4275)(4372) = 42⋅375⋅2 = 46710 Is either of them correct? Explain your reasoning.

OpenStudy (anonymous):

No because 4275 times 4372 is 18690300

OpenStudy (anonymous):

what

OpenStudy (anonymous):

Annabelle and Navene are multiplying (4275)(4372).

OpenStudy (anonymous):

Annabelle's Work (4275)(4372) = 42 + 375 + 2 = 4577

OpenStudy (anonymous):

Navene's Work (4275)(4372) = 42⋅375⋅2 = 46710

OpenStudy (anonymous):

Is either of them correct? Explain your reasoning.

OpenStudy (anonymous):

It's Navene's work because Annabell's work equals 419

OpenStudy (anonymous):

ok thanks \

OpenStudy (anonymous):

also what was navene's answer

OpenStudy (anonymous):

Navene's answer is 31500

OpenStudy (anonymous):

ok thanks